# Solving Fredholm Integral Equations of the Second Kind via Wasserstein Gradient Flows

### v1

@@ 5.1 Simple Gaussian Example
We start by considering a simple integral equation in which the solution $\pi$ can be computed analytically. Consider (1) with $\varphi(x)=(1-\lambda)\mathcal{N}\left(x;0,1\right)$ and $\mathrm{k}(x,y)=\mathcal{N}\left(y;xe^{-\beta},(1-e^{-2\beta})\right)$ for some $10<\lambda<$ 1, $\beta>0$, with $\mathcal{N}(x;\mu,\sigma^{2})$ denoting the density of a normal distribution of mean $\mu$ and variance $\sigma^{2}$ evaluated at $x$. Since $\max_{x}\int_{\mathbb{R}^{d}}\lambda k(x,y)\mathrm{d}y=\lambda<1$, [38, Corollary 2.16] guarantees that (1) admits a unique solution given by $\pi(x)=\mathcal{N}\left(x;0,1\right)$ for all $10<\lambda<$ 1. In our experiments we set $\beta=0.5$.
@@ 5.1.1 Effect of Reference Measure
As discussed in Section 4.2, the reference measure $\pi_{0}$ can be used to impose certain properties on the regularized solution $\pi_{\alpha,\eta}^{\star}$. To check the importance of using a reference measure, we compare the approximate solutions obtained with an improper reference measure (i.e. one corresponding to the entropic penalty in (6)) and several choices of $\pi_{0}$: the target, $\pi_{0}=\pi$, a more diffuse reference measure, $\pi_{0}=\mathcal{N}(0,2^{2})$, and a more concentrated one, $\pi_{0}=\mathcal{N}(0,0.1^{2})$. We set $N=100$, $\gamma=10^{-2}$ and iterate for $n_{T}=200$, which empirically seems sufficient to obtain convergence of the value of $\mathcal{F}_{\alpha}^{\eta}$ (approximated numerically as described in Section 4.2). We consider $\alpha\in[0,1]$.
@@
Figure 1 shows the $\operatorname{ISE}$ and the mean squared error ($\operatorname{MSE}$) for mean and variance of the reconstructions as the value of the regularizing parameter $\alpha$ increases. Large values of $\alpha$ inflate the diffusion coefficient of the MKVSDEs (9) and (15), but, in the latter, there is no $\nabla U$ term in the drift and thus the diffusion term becomes stronger than the drift term for large $\alpha$s, causing the reconstructed solutions to have larger variance than the solution. A reference measure more diffuse than the target has a similar effect.
@@ 5.1.2 Comparison with Reversible jump MCMC
The previous experiments show that the presence of the reference measure significantly improves the results provided by FE2kind-WGF. We now compare our approach with the trans-dimensional MCMC algorithm proposed in [21] which provides a Monte Carlo approximation of the von Neumann series representation of the solution $\pi$ (see, e.g., [38, Section 2.4]). We implement their reversible jump MCMC (RJ-MCMC) with the following set up: we set the death and birth probabilities to 1/3 and use a random walk proposal with variance $\sigma^{2}=0.1^{2}$ for the update move and sample from $\pi$ for the birth move.
For FE2kind-WGF we set $\alpha=0.01$, $N=100$ and $\gamma=10^{-2}$ and iterate for $n_{T}=200$ time steps. We compare three reference measures, the target, $\pi_{0}=\pi$, a more diffuse reference measure, $\pi_{0}=\mathcal{N}(0,2^{2})$, and a more concentrated one, $\pi_{0}=\mathcal{N}(0,0.1^{2})$. We do not consider the case in which no reference measure is given further as the experiments in the previous section showed that this approach does not perform well. The initial distribution of the particles is a Gaussian with mean 0 and small variance ($\sigma^{2}=0.1^{2}$).
We compare the behaviour of the two approaches as $\lambda$ varies in $(0,1)$. We set the number of sample paths drawn using RJ-MCMC to $M=2\cdot 10^{4}$ to match the average cost of FE2kind-WGF (roughly 0.25 seconds for both algorithms). A smooth reconstruction is obtained by using (14) for FE2kind-WGF and using an equivalent strategy for the RJ-MCMC estimator (see [21, Eq. (44)]).
@@ 5.2 Karhunen–Loève Expansions
The Karhunen–Loève expansion of a stochastic process with covariance function $\mathrm{k}(x,y)$ is given by the pairs of eigenvalues and eigenfunctions $(\lambda,\pi)$ of $\mathrm{k}$. For each eigenvalue $\lambda$, the corresponding eigenfunction can be found by solving (1) with $\varphi\equiv 0$. We consider two examples commonly used in the Gaussian process literature [56, Section 4.2]: the squared exponential kernel, $\mathrm{k}(x,y)=\exp(-(y-x)^{2})$, which satisfies our assumptions, and the exponential kernel, $\mathrm{k}(x,y)=\exp(-|y-x|)$, which is not differentiable for $y=x$ and therefore does not satisfy Assumption 1.
The Karhunen–Loève expansion of the exponential kernel is known analytically [27, Section 2.3.3]; the largest eigenvalue is $\lambda=2/(1+\omega^{2})$, where $\omega$ is the largest positive root of $f(\omega)=1-\omega\tan(\omega)$, with corresponding eigenfunction $\pi(x)=\cos(\lambda x)/\sqrt{1+\sin(2\lambda)/(2\lambda)}$ over $[-1,1]$.
We compare the results obtained with the Nyström method, i.e. solving the eigenvalue problem associated with the matrix obtained by discretizing $\mathrm{k}$ over the interval $[-1,1]$, with the results obtained using FE2kind-WGF. We compare the results of the two methods for increasing precision, corresponding to the number of particles $N$ for FE2kind-WGF and the number of discretization intervals for the Nyström method. In particular, we consider $N$ and the number of discretization intervals between 50 and $10^{3}$; for FE2kind-WGF we set $\gamma=1/N$, $\alpha=10^{-2}$ and iterate for $n_{T}=400$. The initial distribution and the reference measure $\pi_{0}$ are both Gaussians centred at 0 and with small variance ($\sigma^{2}=0.05^{2}$).
@@
[…]
@@ 5.3 Equilibrium Distribution of Gaussian Process State Space Models
[7] analyses the equilibrium distribution of Gaussian process state space models (GP-SSM) for a one-dimensional SSM with transition function $f(x)=0.01x^{3}-0.2x^{2}+0.2x$. We fit a GP to learn $f$ using $m$ training pairs, $D_{m}:=(x_{i},f(x_{i})+\epsilon_{i})_{i=1}^{m}$, where $\epsilon_{i}$ are independent Gaussians with mean 0 and standard deviation 5. We follow [7] and use $m=20$ input points $(x_{i})_{i=1}^{m}$ uniformly distributed in $[-5,5]$, and use a squared exponential covariance function $c(x,x^{\prime})=\sigma_{f}^{2}\exp\left(-\left\|x-x^{\prime}\right\|^{2}/(2\ell^{2})\right)$, with $\ell^{2}=3.59^{2}$ and $\sigma_{f}^{2}=4.21^{2}$.
@@
For the Nyström method we use 500 nodes in $[-20,10]$. For FE2kind-WGF we use $N=200,\gamma=0.005$ and iterate for $n_{T}=100$ iterations. The reference measure is a Gaussian with mean 0 and standard deviation 1, we set $\alpha=0.001$. Figure 4 shows the predictive distributions obtained with the two methods. To assess the quality of the results obtained we sample from the predictive distributions $\pi$ obtained with the two algorithms $n=2\cdot 10^{4}$ states $X_{1},\dots,X_{n}$ and apply the transformation $Y_{i}=\mu(X_{i},D_{m})+\sigma(X_{i},D_{m})\xi$, where $\xi$ is a standard Gaussian random variable. To verify that the obtained $\hat{\pi}$ is indeed an invariant distribution we compare our approximations of $\pi$ with the histogram of the $Y_{i}$s, the fit is good for both algorithms (right panel).
@@ 6 Discussion
In this paper we extended the approach of [16] to Fredholm integral equations of the second kind. Under similar assumptions to those in [16] on the kernel $\mathrm{k}$, we show that the regularized functional (4) admits a unique minimizer which is the limiting distribution of the McKean–Vlasov SDE (9).
@@
We show that our method is able to solve a wide variety of equations of the second kind and is competitive with standard methods based on deterministic discretizations. As in the case of [16], the presence of the reference measure $\pi_{0}$ proves beneficial when the problem difficulty increases (Section 5.1). In addition, FE2kind-WGF is robust to small deviations from Assumption 1, as shown in Section 5.2 where we consider a kernel $\mathrm{k}$ which is not continuously differentiable.
While we focused on one dimensional examples to benchmark our method against other algorithms proposed in the literature, we believe that FE2kind-WGF has the potential to tackle higher dimensional problems and, as shown in [16] for equations of the first kind, to significantly outperform methods based on deterministic discretizations. One area in which this could prove particularly beneficial is spatial statistics, in which multidimensional spatial Karhunen–Loève expansions are often employed [26, 15].
@@ Acknowledgments
AMJ acknowledges financial support from the Engineering and Physical Sciences Research Council (EPSRC; grants EP/R034710/1 and EP/T004134/1) and by United Kingdom Research and Innovation (UKRI) via grant EP/Y014650/1, as part of the ERC Synergy project OCEAN. For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising from this submission.
Data access statement: No new data was created during this research. Julia code to reproduce all examples is available online at https://github.com/FrancescaCrucinio/FE2kind_WGF.
@@ Appendix A Proof of Proposition 1
Before proceeding to the proof of Proposition 1 we state and prove the following auxiliary result which is adapted from [25, Appendix A.1].
@@
The KL divergence above can be written as an $f$-divergence for $f(s)=s\log s$ with the convention $00\log 0=$ 0, $\textup{KL}\left(\mu\middle|\nu\right)=\int\limits_{\mathbb{R}^{d}}f(\mathrm{d}\mu/\mathrm{d}\nu)\mathrm{d}\nu$. The convex conjugate of $f$ is, for all $u\in\mathbb{R}$,
@@
    by definition since both $\pi$ and $\nu(\mathrm{d}x)=\left(\varphi(x)+\lambda\int\limits_{\mathbb{R}^{d}}\mathrm{k}(x,y)\pi(y)\mathrm{d}y\right)\mathrm{d}x$ are probability measures. To show that $\mathcal{F}(\pi)$ is convex, consider the functional $\mathcal{M}:\pi\mapsto\left(\varphi+\lambda\int\limits_{\mathbb{R}^{d}}\mathrm{k}(\cdot,y)\mathrm{d}\pi(y)\right)$ which is linear w.r.t $\pi$ and therefore convex. Thus, we have that $\mathcal{F}(\pi)$ can be written as $\textup{KL}\left(\pi\middle|\mathcal{M}(\pi)\right)$. In addition, [22, Lemma 1.4.3-(b)] guarantees that the KL between probability distributions is jointly convex in both arguments. It follows that, for all $t\in[0,1]$
@@
    and take $f(s)=s\log s$ with the convention $00\log 0=$ 0. Then, $\mathcal{F}^{\eta}(\pi)$ can be written as $\int\limits_{\mathbb{R}^{d}}f(\mathrm{d}\pi/\mathrm{d}\nu)\mathrm{d}\nu$ with $\nu(\mathrm{d}x)=\left(\varphi(x)+\lambda\int\limits_{\mathbb{R}^{d}}\mathrm{k}(x,y)\pi(y)\mathrm{d}y+\eta\right)\mathrm{d}x$ a $\sigma$-finite measure, while $\pi$ is a probability measure. Using Jensen’s inequality on the convex function $f$ w.r.t. the probability measure $\nu-\eta$ we have
    $\int\limits_{\mathbb{R}^{d}}f\left(\frac{\mathrm{d}\pi}{\mathrm{d}\nu}(x)\right)\nu(\mathrm{d}x)$
    $\qquad=\int\limits_{\mathbb{R}^{d}}f\left(\frac{\mathrm{d}\pi}{\mathrm{d}\nu}(x)\right)(\nu(\mathrm{d}x)-\eta\mathrm{d}x)+\eta\int\limits_{\mathbb{R}^{d}}f\left(\frac{\mathrm{d}\pi}{\mathrm{d}\nu}(x)\right)\mathrm{d}x$
    $\qquad\geq f\left(\int\limits_{\mathbb{R}^{d}}\frac{\mathrm{d}\pi}{\mathrm{d}\nu}(x)(\nu(\mathrm{d}x)-\eta\mathrm{d}x)\right)+\eta\int\limits_{\mathbb{R}^{d}}f\left(\frac{\mathrm{d}\pi}{\mathrm{d}\nu}(x)\right)\mathrm{d}x.$
@@
    where we used Assumption 1 to obtain the lower bound. Using (19,20) and the definition of $\mathcal{F}^{\eta}$ we find
    $\mathcal{F}^{\eta}(\pi)\geq$ $-e^{-1}+\frac{\eta}{(1+\lambda)\mathtt{M}+\eta}\mathcal{F}^{\eta}(\pi)$
    $\left(1-\frac{\eta}{(1+\lambda)\mathtt{M}+\eta}\right)\mathcal{F}^{\eta}(\pi)\geq$ $-e^{-1}$
    $\mathcal{F}^{\eta}(\pi)\geq$ $-\left(1-\frac{\eta}{(1+\lambda)\mathtt{M}+\eta}\right)^{-1}e^{-1}:=\mathtt{C}$
    The convexity of $\mathcal{F}^{\eta}$ follows from Lemma 1 with $\mu=\pi$ and $\nu(\mathrm{d}x)=\left(\varphi(x)+\lambda\int\limits_{\mathbb{R}^{d}}\mathrm{k}(x,y)\pi(y)\mathrm{d}y+\eta\right)\mathrm{d}x$ since the fact that $\eta>0$ guarantees that any probability measure with a Lebesgue density is absolutely continuous w.r.t. $\nu$.
@@
    To see that $\mathcal{F}_{\alpha}^{\eta}$ is proper observe that, since $\textup{KL}\left(\pi\middle|\pi_{0}\right)\geq 0$, $\mathcal{F}_{\alpha}^{\eta}(\pi)\geq\mathtt{C}$ for all $\pi\in\mathcal{P}(\mathbb{R}^{d})$.
    Now take a reference measure $\pi_{0}$ with finite entropy, i.e. $|\mathrm{H}(\pi_{0})|<+\infty$. We have that
@@
    Hence, there exists $\pi=\pi_{0}$ for which $\mathcal{F}_{\alpha}^{\eta}(\pi)<+\infty$, showing that the functional is proper.
    Let $(\pi_{n})_{n\geq 1}\in(\mathcal{P}(\mathbb{R}^{d}))^{\mathbb{N}}$ be such that $\lim_{n\to+\infty}\pi_{n}=\pi\in\mathcal{P}(\mathbb{R}^{d})$. Since for any $(x,y)\in\mathbb{R}^{d}\times\mathbb{R}^{d}$, $|\mathrm{k}(x,y)|\leq\mathtt{M}$ and $\mathrm{k}$ is continuous, we have that $\lim_{n\to+\infty}\pi_{n}[\mathrm{k}(x,\cdot)]=\pi[\mathrm{k}(x,\cdot)]$ for each $x$. This and the fact that the Kullback–Leibler divergence is lower semi-continuous in both arguments [22, Lemma 1.4.3-(b)] guarantees that $\mathcal{F}^{\eta}$ is lower semi-continuous. [22, Lemma 1.4.3-(b)] also guarantees that $\textup{KL}\left(\pi\middle|\pi_{0}\right)$ is strictly convex and lower semi-continuous. It follows that $\mathcal{F}^{\eta}_{\alpha}$ is strictly convex and lower semi-continuous.
    To see that $\mathcal{F}^{\eta}_{\alpha}$ is coercive observe that $\mathcal{F}^{\eta}_{\alpha}$ is the sum of the lower bounded lower semi-continuous functional $\mathcal{F}^{\eta}$ and the coercive functional [22, Lemma 1.4.3-(c)] $\textup{KL}\left(\pi\middle|\pi_{0}\right)$, then for any $\beta\in\mathbb{R}$
    $S$ :=$:=\left\{\pi\in\mathcal{P}(\mathbb{R}^{d}):\mathcal{F}^{\eta}(\pi)+\alpha\textup{KL}\left(\pi\middle|\pi_{0}\right)\leq\beta\right\}$
    $\subseteq\left\{\pi\in\mathcal{P}(\mathbb{R}^{d}):\alpha\textup{KL}\left(\pi\middle|\pi_{0}\right)\leq\beta-\mathtt{C}\right\}:=\tilde{S},$

### v2

@@ 5.1 Simple Gaussian Example
We start by considering a simple analytically-tractable integral equation. Consider, for $\lambda\in(0,1)$,  (1) with $\varphi(x)=(1-\lambda)\mathcal{N}\left(x;0,1\right)$ and $\mathrm{k}(x,y)=\mathcal{N}\left(y;xe^{-\beta},1-e^{-2\beta}\right)$ for some $\beta>0$, with $\mathcal{N}(x;\mu,\sigma^{2})$ denoting the density of a normal distribution of mean $\mu$ and variance $\sigma^{2}$ evaluated at $x$. Since $\max_{x}\int_{\mathbb{R}^{d}}\lambda k(x,y)\mathrm{d}y=\lambda<1$, [40, Cor. 2.16] guarantees that (1) admits a unique solution, $\pi(x)=\mathcal{N}\left(x;0,1\right)$ for all $0<\lambda<1$. In our experiments we set $\beta=0.5$.
@@ 5.1.1 Effect of Reference Measure
As discussed in Section 4.2, the reference measure $\pi_{0}$ can be used to impose certain properties on the regularized solution $\pi_{\alpha,\eta}^{\star}$. To better understand the role of the reference measure, we compare the approximate solutions obtained with an improper reference measure (i.e. one corresponding to the entropic penalty in (6)) and several choices of $\pi_{0}$: the target, $\pi_{0}=\pi$, a more diffuse reference measure, $\pi_{0}=\mathcal{N}(0,2^{2})$, and a more concentrated one, $\pi_{0}=\mathcal{N}(0,0.1^{2})$. We set $N=100$ and $\gamma=10^{-2}$ as suggested in Section 4.2 and iterate for $n_{T}=200$, which empirically seems sufficient to obtain convergence of the value of $\mathcal{F}_{\alpha}^{\eta}$ (approximated numerically as described in Section 4.2). We consider $\alpha\in[0,1]$.
@@
[…]
@@ 5.1.2 Comparison with Reversible jump MCMC
The previous experiments show that the presence of the reference measure significantly improves the results provided by FE2kind-WGF. We now compare our approach with the MCMC algorithm of [23] which provides a Monte Carlo approximation of the von Neumann series representation of the solution $\pi$ (see, e.g., [40, Sec. 2.4]). We implement their reversible jump MCMC (RJ-MCMC) with the following set up: we set the death and birth probabilities to 1/3 and use a random walk proposal with variance $\sigma^{2}=0.1^{2}$ for the update move and sample from $\pi$ for the birth move.
For FE2kind-WGF we set $\alpha=0.01$, $N=100$ and $\gamma=10^{-2}$ and iterate for $n_{T}=200$ time steps which empirically seems sufficient to obtain convergence of the value of $\mathcal{F}_{\alpha}^{\eta}$ (approximated numerically as described in Section 4.2). We compare three reference measures, the target, $\pi_{0}=\pi$, a more diffuse reference measure, $\pi_{0}=\mathcal{N}(0,2^{2})$, and a more concentrated one, $\pi_{0}=\mathcal{N}(0,0.1^{2})$. We do not consider the case in which no reference measure is given further as the experiments in the previous section showed that this approach does not perform well. The initial distribution of the particles is a Gaussian with mean 0 and small variance ($\sigma^{2}=0.1^{2}$).
We compare the behaviour of the two approaches as $\lambda$ varies in $(0,1)$. We set the number of sample paths drawn using RJ-MCMC to $M=2\cdot 10^{4}$ to match the average cost of FE2kind-WGF (roughly 0.25 seconds for both algorithms). A smooth reconstruction is obtained by using (15) for FE2kind-WGF and using an equivalent strategy for the RJ-MCMC estimator (see [23, Eq. (44)]).
@@ 5.2 Karhunen–Loève Expansions
The Karhunen–Loève expansion of a given stochastic process with covariance function $\mathrm{k}(x,y)$ is given by the pairs of eigenvalues and eigenfunctions $(\lambda,\pi)$ of $\mathrm{k}$. For any eigenvalue $\lambda$, the corresponding eigenfunction can be found by solving (1) with $\varphi\equiv 0$. We consider two examples commonly used in the Gaussian process literature [60, Sec. 4.2]: the squared exponential kernel, $\mathrm{k}(x,y)=\exp(-(y-x)^{2})$, which satisfies our assumptions, and the exponential kernel, $\mathrm{k}(x,y)=\exp(-|y-x|)$, which is not differentiable for $y=x$ and therefore does not satisfy Assumption 1.
The Karhunen–Loève expansion of the exponential kernel is known analytically [29, Sec. 2.3.3]; the largest eigenvalue is $\lambda=2/(1+\omega^{2})$, where $\omega$ is the largest positive root of $f(\omega)=1-\omega\tan(\omega)$, with corresponding eigenfunction $\pi(x)=\cos(\lambda x)/\sqrt{1+\sin(2\lambda)/(2\lambda)}$ over [1,1][-1,1].
We compare the results obtained with the Nyström method, i.e. solving the eigenvalue problem associated with the matrix obtained by discretizing $\mathrm{k}$ over the interval [1,1][-1,1], with the results obtained using FE2kind-WGF. We compare the results of the two methods for increasing precision, corresponding to the number of particles $N$ for FE2kind-WGF and the number of discretization intervals for the Nyström method. In particular, we consider $N$ and the number of discretization intervals between 50 and $10^{3}$; for FE2kind-WGF we set $\gamma=1/N$ as suggested in Section 4.2, $\alpha=10^{-2}$ and iterate for $n_{T}=400$ which empirically seems sufficient to obtain convergence of the value of $\mathcal{F}_{\alpha}^{\eta}$ (approximated numerically as described in Section 4.2). The initial distribution and the reference measure $\pi_{0}$ are both Gaussians centred at 0 and with small variance ($\sigma^{2}=0.05^{2}$), in this case the reference measure carries information on the support of $\pi$, whose mass is mostly concentrated on [1,1][-1,1].
@@
Refer to captionNNGain in $\operatorname{ISE}$Refer to captionxxRefer to captionxxsquared exponentialexponential
@@ 5.3 Equilibrium Distribution of Gaussian Process State Space Models
[7] analyses the equilibrium distribution of Gaussian process state space models (GP-SSM) for a one-dimensional SSM with transition function $f(x)=0.01x^{3}-0.2x^{2}+0.2x$. We fit a GP to learn $f$ using $m$ training pairs, $D_{m}:=(x_{i},f(x_{i})+\epsilon_{i})_{i=1}^{m}$, where $\epsilon_{i}$ are independent Gaussians with mean 0 and standard deviation 5. We follow [7] and use $m=20$ input points $(x_{i})_{i=1}^{m}$ uniformly distributed in [5,5][-5,5], and use a squared exponential covariance function $c(x,x^{\prime})=\sigma_{f}^{2}\exp\left(-\left\|x-x^{\prime}\right\|^{2}/(2\ell^{2})\right)$, with $\ell^{2}=3.59^{2}$ and $\sigma_{f}^{2}=4.21^{2}$.
@@
For the Nyström method we use 500 nodes in [20,10][-20,10]. For FE2kind-WGF we use $N=200$ and $\gamma=1/N$ as suggested in Section 4.2 and iterate for $n_{T}=100$ iterations which empirically seems sufficient to obtain convergence of the value of $\mathcal{F}_{\alpha}^{\eta}$ (approximated numerically as described in Section 4.2). The reference measure is a Gaussian with mean 0 and standard deviation 1 (this is motivated by the fact that $\pi$ assigns positive mass to 0 and our empirical studies show that concentrated distributions tend to lead to faster convergence than diffuse ones), we set $\alpha=0.001$. Figure 4 shows the predictive distributions obtained with the two methods. To assess the quality of the results obtained we sample from the predictive distributions $\pi$ obtained with the two algorithms $n=2\cdot 10^{4}$ states $X_{1},\dots,X_{n}$ and apply the transformation $Y_{i}=\mu(X_{i},D_{m})+\sigma(X_{i},D_{m})\xi$, where $\xi$ is a standard Gaussian random variable. To verify that the obtained $\hat{\pi}$ is indeed an invariant distribution we compare our approximations of $\pi$ with the histogram of the $Y_{i}$s, the fit is good for both algorithms (right panel).
@@ 6 Discussion
In this paper we extended the approach of [16] to Fredholm integral equations of the second kind. Under similar assumptions to those in [16] on the kernel $\mathrm{k}$, we show that the regularized functional (4) admits a unique minimizer which is the limiting distribution of the McKean–Vlasov SDE (9). In addition, we study the limiting behaviour of the minimizer when the regularization parameter tends to 0 (Proposition 2).
@@
Our method addresses a wide variety of equations of the second kind and is competitive with standard methods based on deterministic discretizations. As in the case of [16], the presence of the reference measure $\pi_{0}$ proves beneficial when the problem difficulty increases (Section 5.1). In addition, FE2kind-WGF is robust to small deviations from Assumption 1, as shown in Section 5.2 where we consider a kernel $\mathrm{k}$ which is not continuously differentiable.
While we focused on one dimensional examples to benchmark our method against other algorithms proposed in the literature, we believe that FE2kind-WGF has the potential to tackle higher dimensional problems and, as shown in [16, Section 5.4] for equations of the first kind, to significantly outperform methods based on deterministic discretizations with the same storage cost. This is also confirmed in the one-dimensional case by Figure 3 left panel. One area in which this could prove particularly beneficial is spatial statistics, in which multidimensional spatial Karhunen–Loève expansions are often employed [28, 15].
@@ Acknowledgments
AMJ acknowledges the financial support of the United Kingdom Engineering and Physical Sciences Research Council (EPSRC; grants EP/R034710/1 and EP/T004134/1) and by United Kingdom Research and Innovation (UKRI) via grant number EP/Y014650/1, as part of the ERC Synergy project OCEAN.
FRC gratefully acknowledges the “de Castro” Statistics Initative at the Collegio Carlo Alberto and the Fondazione Franca e Diego de Castro. FRC is supported by the Gruppo Nazionale per l’Analisi Matematica, la Probabilità e le loro Applicazioni (GNAMPA-INdAM).
For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising from this submission.
Data access statement: No new data was created during this research. Julia code to reproduce all examples is available at
https://github.com/FrancescaCrucinio/FE2kind_WGF.
@@ Appendix A Proof of Proposition 1
A.1 Auxiliary Result for the Proof of Proposition 1
The following auxiliary result is adapted from [27, Appendix A.1].
@@ A.1 Auxiliary Result for the Proof of Proposition 1
The KL divergence above can be written as an $f$-divergence for $f(s)=s\log s$ with the convention $0\log 0=0$, $\textup{KL}\left(\mu\middle|\nu\right)=\int\limits_{\mathbb{R}^{d}}f(\mathrm{d}\mu/\mathrm{d}\nu)\mathrm{d}\nu$. The convex conjugate of $f$ is, for all $u\in\mathbb{R}$,
@@ A.2 Proof of Proposition 1
    because $\pi$ and $\nu(\mathrm{d}x)=\left(\varphi(x)+\lambda\int\limits_{\mathbb{R}^{d}}\mathrm{k}(x,y)\pi(y)\mathrm{d}y\right)\mathrm{d}x$ are probabilities. To show that $\mathcal{F}(\pi)$ is convex, consider the functional
    $\mathcal{M}:\pi\mapsto\left(\varphi+\lambda\int\limits_{\mathbb{R}^{d}}\mathrm{k}(\cdot,y)\mathrm{d}\pi(y)\right)$
    which is linear w.r.t $\pi$ and therefore convex. Thus, we have that $\mathcal{F}(\pi)$ can be written as $\textup{KL}\left(\pi\middle|\mathcal{M}(\pi)\right)$. In addition, [24, Lemma 1.4.3-(b)] guarantees that the KL between probability distributions is jointly convex in both arguments. Thus, for all $t\in[0,1]$
@@
    and take $f(s)=s\log s$ with the convention $0\log 0=0$. Then, $\mathcal{F}^{\eta}(\pi)$ can be written as $\int\limits_{\mathbb{R}^{d}}f(\mathrm{d}\pi/\mathrm{d}\nu)\mathrm{d}\nu$ with $\nu(\mathrm{d}x)=\left(\varphi(x)+\lambda\int\limits_{\mathbb{R}^{d}}\mathrm{k}(x,y)\pi(y)\mathrm{d}y+\eta\right)\mathrm{d}x$ a $\sigma$-finite measure, while $\pi$ is a probability measure. Using Jensen’s inequality on the convex function $f$ w.r.t. the probability measure $\nu-\eta$ we have
    $\int\limits_{\mathbb{R}^{d}}f\left(\frac{\mathrm{d}\pi}{\mathrm{d}\nu}(x)\right)\nu(\mathrm{d}x)$ $=\int\limits_{\mathbb{R}^{d}}f\left(\frac{\mathrm{d}\pi}{\mathrm{d}\nu}(x)\right)(\nu(\mathrm{d}x)-\eta\mathrm{d}x)+\eta\int\limits_{\mathbb{R}^{d}}f\left(\frac{\mathrm{d}\pi}{\mathrm{d}\nu}(x)\right)\mathrm{d}x$
    $\geq f\left(\int\limits_{\mathbb{R}^{d}}\frac{\mathrm{d}\pi}{\mathrm{d}\nu}(x)(\nu(\mathrm{d}x)-\eta\mathrm{d}x)\right)+\eta\int\limits_{\mathbb{R}^{d}}f\left(\frac{\mathrm{d}\pi}{\mathrm{d}\nu}(x)\right)\mathrm{d}x.$
@@
    where we used Assumption 1 to obtain the lower bound. Using (20,21) and the definition of $\mathcal{F}^{\eta}$ we find
    $\mathcal{F}^{\eta}(\pi)\geq$ $-e^{-1}+\frac{\eta}{(1+\lambda)\mathtt{M}+\eta}\mathcal{F}^{\eta}(\pi)\geq-\left(1-\frac{\eta}{(1+\lambda)\mathtt{M}+\eta}\right)^{-1}e^{-1}=:\mathtt{C}$
    The fact that $\eta>0$ guarantees that any probability measure with a Lebesgue density is absolutely continuous w.r.t. $\nu$ and so the convexity of $\mathcal{F}^{\eta}$ follows from Lemma 1 with $\mu=\pi$ and
    $\nu(\mathrm{d}x)=\left(\varphi(x)+\lambda\int\limits_{\mathbb{R}^{d}}\mathrm{k}(x,y)\pi(y)\mathrm{d}y+\eta\right)\mathrm{d}x.$
@@
    To see that $\mathcal{F}_{\alpha}^{\eta}$ is proper observe that, since $\textup{KL}\left(\pi\middle|\pi_{0}\right)\geq 0$, $\mathcal{F}_{\alpha}^{\eta}(\pi)\geq\mathtt{C}$ for all $\pi\in\mathcal{P}(\mathbb{R}^{d})$. Take a reference measure $\pi_{0}$ with finite entropy, i.e. $|\mathrm{H}(\pi_{0})|<+\infty$. We have
@@
    Hence, there exists $\pi=\pi_{0}$ for which $\mathcal{F}_{\alpha}^{\eta}(\pi)<+\infty$ and the functional is proper.
    Let $(\pi_{n})_{n\geq 1}\in(\mathcal{P}(\mathbb{R}^{d}))^{\mathbb{N}}$ be such that $\lim_{n\to+\infty}\pi_{n}=\pi\in\mathcal{P}(\mathbb{R}^{d})$. Since for any $(x,y)\in\mathbb{R}^{d}\times\mathbb{R}^{d}$, $|\mathrm{k}(x,y)|\leq\mathtt{M}$ and $\mathrm{k}$ is continuous, we have that $\lim_{n\to+\infty}\pi_{n}[\mathrm{k}(x,\cdot)]=\pi[\mathrm{k}(x,\cdot)]$ for each $x$. This and the fact that the Kullback–Leibler divergence is lower semi-continuous in both arguments [24, Lemma 1.4.3-(b)] guarantees that $\mathcal{F}^{\eta}$ is lower semi-continuous. [24, Lemma 1.4.3-(b)] also guarantees that $\textup{KL}\left(\pi\middle|\pi_{0}\right)$ is strictly convex and lower semi-continuous. It follows that $\mathcal{F}^{\eta}_{\alpha}$ is strictly convex and lower semi-continuous.
    To see that $\mathcal{F}^{\eta}_{\alpha}$ is coercive observe that $\mathcal{F}^{\eta}_{\alpha}$ is the sum of the lower bounded lower semi-continuous functional $\mathcal{F}^{\eta}$ and the coercive functional [24, Lemma 1.4.3-(c)] $\textup{KL}\left(\pi\middle|\pi_{0}\right)$, then for any $\beta\in\mathbb{R}$
    $S:=$ $\left\{\pi\in\mathcal{P}(\mathbb{R}^{d}):\mathcal{F}^{\eta}(\pi)+\alpha\textup{KL}\left(\pi\middle|\pi_{0}\right)\leq\beta\right\}$
    $\subseteq$ $\left\{\pi\in\mathcal{P}(\mathbb{R}^{d}):\alpha\textup{KL}\left(\pi\middle|\pi_{0}\right)\leq\beta-\mathtt{C}\right\}=:\tilde{S}$,